OpenStudy (anonymous):

Find the indefinite integral of sin(x degrees)/4

zepdrix (zepdrix):

\[\large \int\limits \frac{\sin x}{4}\;dx\]This is your integral?

OpenStudy (anonymous):

So by indefinite integral, we're really looking for the anti-derivative. Simply ask yourself what has a derivative of sin(x)?

OpenStudy (anonymous):

Derivative of sin(x) = -cos(x) And that is my integral, but the x has a degree sign above it

OpenStudy (anonymous):

I don't really think that matters. It doesn't matter if you're talking about \[\sin \theta, \sin(x)\]the point is that there's a variable inside.

OpenStudy (anonymous):

Sorry, that was integral

hartnn (hartnn):

it matters like this, those standard formulae are for angles in radians, so you need to convert degrees into radians \(x^o=(x \times \pi/180)^r\)

OpenStudy (anonymous):

what i have at the moment is this: